OpenStudy (anonymous):

Hey everyone, could someone help me on this question? Find the derivative: g(x) = 7/(x^2+2sqrootx)^3 Hey everyone, could someone help me on this question? Find the derivative: g(x) = 7/(x^2+2sqrootx)^3 @Mathematics

OpenStudy (anonymous):

do you mind using the equation tool i dont understand the problem

OpenStudy (turingtest):

convert the denominator to a negative exponent, and the square root to a fractional exponent:\[g(x) ={ 7\over(x^2+2\sqrt{x})^3}=7(x^2+2x^{1/2})^{-3}\] from here on use the chain rule and the problem should be pretty straightforward.

OpenStudy (anonymous):

sure thing, sorry for the confusion, \[7/\left( x^{2} +2\sqrt{x} \right)^{3}\]

OpenStudy (anonymous):

Oh okay then, so I won't have to use the quotient rule just the chain rule.. thanks!